3. Consider a competitive firm with total costs given by TC(q) = 100 + 10q+q? (e) Graph the ATC, AVC, MC, and MR curves in aThe market price is p=50
Answer #1

TC +100+0q+q? AC/ Ria MC/AVC AC= 100+ 10+q AVC AC Arc = lota Mc= 10t2q e) Profit is max when Mrame 10+29=50 19=204 AC=100, 10